OpenStudy (anonymous):

if alpha anb beta be the roots of equation: X2-px+q =0 and if Vn=@n+Bn prove that Vn+1=PVn-Qvn-1

OpenStudy (anonymous):

???

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@dumbcow, can you make heads of tails of this?

OpenStudy (anonymous):

HEY dumo..wht u wana say?

OpenStudy (dumbcow):

so if a,b is alpha and beta then a,b are roots of x^2 -px+q v = a+b prove v+1 = pv -qv -1

OpenStudy (dumbcow):

(x-a)(x-b) = x^2 -px +q ab = q a+b = p --> a+b+1 = (a+b)^2 -ab(a+b) -1 ??

OpenStudy (anonymous):

hmm..